Suppose that an economy is characterized by the following equations:



C = 400 + 0.75(Y - T)



a = 400 - 25r



G = 200



T = 200



to the private



MS = 2, 000 Derive:



i IS curve



P = 2



(M / P) * d = V - 200r



ii LM curve.



iii Equilibrium level of income and interest rate.



b) Kenya has recently lifted the interest rate cap in a move to revive shrinking cost access



sector. Do you agree with the move? Discuss c) Trade war happens when one country retaliates against another by using import tariffs or placing other restriction on the other country's imports. This can commence if one country perceives a competitor nation has unfair trading practices. Explain the policy implications of trade dispute in an economy.




1
Expert's answer
2021-11-28T17:49:04-0500

a)

i. The IS curve is defined by the equation:

Y=C(Y-T(Y))+I(r)+G+NX(Y)

Y=400 + 0.75(Y - T)+ 400 - 25r+ 200

Y=400 + 0.75(Y - 200)+ 400 - 25r+ 200

Y=400+0.75Y-150+400-25r+200

Y=850+0.75Y-25r

Y-0.75Y=800-25r

0.25Y=800-25r

Y=3200-100r


ii. M(s)=M(d)

(M / P) * d = Y- 200r

(M / 2) * d = 3200-100r- 200r

M(d)/2=3200-300r

M(d)=M(s)=6400-600r


iii. 3200-100r=6400-600r

500r=3200

r=6.4

Y=3200-100r=3200-100*6.4=3200-640=2560


b)Agree. The removal of the upper limit of the rate increases the risk of price changes Kenyan banks, expect higher rates of credit growth in the next 12-18 months and an increase in lending to those sectors of the economy that had limited development and access to loans


c)The outcome of a trade war is either the loss of economic status by countries, or their complete ruin (if the war is successfully waged - for competitors, if unsuccessful - for the country that started the trade war, or for both countries).In trade wars, as a rule, there are no winners, the mutual exchange of measures to restrict imports leads to a reduction in the volume of international trade, affecting to one degree or another all its participants and reducing to one degree or another the economic well-being of absolutely all countries involved in the conflict. Sometimes the consequences of the establishment of trade barriers entail very serious consequences.